OpenStudy (anonymous):

Solve the system of equations y= -0.04x^2 + 8.3x , y=1.12x + 1.6

OpenStudy (anonymous):

BY EQUATING THE EQUATIONS -0.04X^2+8.3X=1.12X+1.6 0.04X^2-7.18X+1.6 THEN FACTORIZE

OpenStudy (anonymous):

y= -0.04x^2 + 8.3x -(1) y=1.12x + 1.6 -(2) Put (1) into (2) 1.12x + 1.6 = -0.04x^2 + 8.3x 0.04x^2 - 7.18x + 1.6 =0 4x^2 - 718x + 160 =0 Oh.. then use quadratic formula to solve it!

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt {b^2-4ac}}{2a}\] Put a = 4, b= -718 and c=160 into the above formula. See what you can get!~

OpenStudy (anonymous):

I have fell off the cliff now!@!

OpenStudy (anonymous):

Don't worry, superman will save you :)

OpenStudy (anonymous):

Before superman arrives, let save yourself first :) \[x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]\[x = \frac{-(-718) \pm \sqrt{(-718)^2-4(4)(160)}}{2(4)}\]\[x = \frac{718 \pm \sqrt{512964}}{8} \]\[x= \frac{359 \pm \sqrt{128241}}{4} \]\[=\frac{359 \pm 3\sqrt{14249}}{4} \] It's really ...not elegant..